OpenStudy (anonymous):

simplify 3x+9+15x+45=

OpenStudy (anonymous):

18x + 54=0 -54 -54 18x=-54 /18 /18 x=-3

OpenStudy (texaschic101):

3x + 9 + 15x + 45 (combine like terms) 3x + 15x +45 + 9 18x + 54 I am not sure if you need to factor or not...but it will factor... 18(x + 3)